# Diffusion MRI with Machine Learning

### v1

@@ II-C Tractography and structural connectivity
  • An increasingly popular class of models in tractography is based on recurrent neural networks (RNNs) [150, 147, 148, 152]. Poulin et al. [150] trained MLPs and Gated Recurrent Units (GRUs) [153] to map the dMRI signal to the streamline directions. The input to their model was the b0-normalized signal, resampled to a fixed spherical grid with 100 directions. The information from the preceding tracing steps was passed to the current step via the hidden state vector. Experiments showed that the GRU-based method outperformed 96 competing techniques on a simulated dataset and achieved expert-level results on a scan from the HCP dataset. Benou et al. [148] used an RNN to estimate a probabilistic FOD, which could then be used for deterministic or probabilistic tractography. The streamline tracing was stopped when the model predicted a termination label or when the entropy of the predicted FOD was below a threshold. This method showed competitive performance compared with classical methods and machine learning techniques, especially in terms of false positive rate and ability to reconstruct various fiber bundles.
  • The majority of the proposed methods have followed a supervised learning strategy. The main difficulty here is obtaining rich data with reliable ground truth. One approach to obtaining ground truth tractography data is via elaborate manual editing of automatically generated tractograms by experts [154]. However, even for experts it is often impossible to resolve ambiguous situations and this practice suffers from high inter-observer variability [155, 156]. Physical phantoms represent an alternative [157, 158]. Poupon et al. have developed the FiberCup phantom to simulate a coronal slice of the brain at the level of corticospinal tracts [159, 160]. This phantom includes a rich set of crossing, fanning, and kissing tracts as well as U-fibers. The main problem with this approach is that no phantom can represent the complexity and inter-subject variability of the human brain.
@@
  • Given the difficulty of obtaining training data, there is a growing interest in reinforcement learning techniques [161, 151, 162]. Theberge et al. developed the first deep reinforcement learning-based tractography method [151]. In their method, the state is derived from the diffusion signal while the reward is computed from the FOD peaks and the direction of the streamline. The model input consists of SH-represented FOD at the current voxel and its six neighbors. Additionally, the past four streamline directions are included as extra inputs to the model. The reward function is based on local information. Specifically, it promotes the closeness of the predicted streamline tracing step with the major FOD peaks and encourages streamline smoothness. Experiments show that the reinforcement learning method is competitive with supervised learning methods while also having better generalizability to unseen datasets. However, a true reinforcement learning method requires the knowledge of streamline start and end points, which is unavailable for in-vivo data.
@@
  • In an interesting work, Cai et al. proposed to learn a tractography model based on anatomical (T1w) MRI [44]. Their model, based on convolutional and recurrent neural networks, was trained using a teacher-student approach. The teacher model was trained to predict the streamlines based on the dMRI signal. The student model was then trained by the teacher model to make the same prediction based on the T1w image. Experiments showed that the accuracy of tractography and structural connectivity measures from this method were similar to scan-rescan variability of standard dMRI-based results.
II-D Delineation of white matter tracts
@@
The brain white matter is organized into distinct tracts, which consist of bundles of myelinated axons that connect different brain regions such as the cerebral cortex and the deep gray matter [164, 165, 166]. Although they are tightly packed and often cross one another, each tract has a different function and connects different regions of the brain. Accurate segmentation of these tracts is needed in clinical studies and medical research. For example, in surgical planning one needs to know the precise extent of the individual tracts in order to assess the risk of damage to specific neurocognitive functions that may result from surgical removal of brain tissue [167, 139]. Furthermore, changes in the tissue micro-structure on specific tracts can be associated with brain development and disorders [168, 169, 170].
@@
Diffusion MRI is the modality of choice for non-invasive delineation of white matter tracts [166]. Individual tracts may be extracted from whole-brain tractograms by specifying inclusion and exclusion regions of interest (ROIs). This process, which is usually referred to as “virtual dissection”, is time-consuming, requires substantial expertise, and suffers from high inter-observer variability and low reproducibility [156, 171]. Some prior works have aimed at automating the virtual dissection process by learning to compute the inclusion/exclusion ROIs [172, 173, 174]. Moreover, there has been much effort to develop fully automatic methods for tract delineation. These methods can be divided into (1) methods that start by performing streamline tractography and (2) methods that do not depend on tractography. We have summarized some of the works in these two categories of methods, respectively, in Tables VI and VII. Given the large number of published methods, these tables only list a selection of more recent methods with a focus on deep learning techniques. Below, we explain our main findings from our study of these methods.
@@
TABLE VI: A summary of selected tractography-based methods for automatic delineation/segmentation of white matter tracts.
 Reference
Model
Input
Number of tracts
Main findings
 [175]
CNN
tractogram
54
@@
Superior accuracy across the human lifespan from neonatal to old age. The method works well on brains with gross pathologies such as tumors.
[176]
graph CNN
tractogram
12
Superior accuracy compared with standard methods, especially for small tracts such as fornix and commissure anterior.
[177]
unclear
tractogram
1
Only addresses the arcuate fasciculus.
[178]
bidirectional LSTM
tractogram
8
The method achieves streamline classification accuracy of $>96\%$ and recall of $>73\%$.
[179, 180]
CNN
tractogram
17
Low false positive rate compared with competing methods.
[181]
CNN
tractogram
64
The focus of this method is on detecting eloquent white matter tracts for epilepsy surgery.
[182]
spectral embedding clustering
tractogram
10
@@
The method achieves high tract segmentation accuracy and can be useful for cross-subject white matter studies.
[183]
restricted Boltzmann machines (RBM)
tractogram
61
@@
The method offers memory-efficient and unsupervised detection of anatomically relevant tracts without the need to explicitly define the tracts.
[173]
Bayesian model
tractogram and anatomical MRI
18
@@
The model trained on healthy subjects worked well on schizophrenia patients.
[184]
LSTM
tractogram
8
The proposed method is particularly useful in situations when the test brains are arbitrarily oriented.
[185]
Siamese Graph convolutional neural network
tractogram and gray matter parcellation
not clear
@@
The method produces accurate delineation of tracts across age and gender for healthy as well as diseased brains.
[186]
CNN
orientation of the major FOD peaks
72
This work showed that modeling the label correlations improved the segmentation accuracy, especially for tracts that were more difficult to segment.
[187]
MLP
tractogram
198
Point cloud-based representation of the streamlines and contrastive supervised learning enables accurate extraction of superficial white matter tracts.
[188]
normalized cuts
tractogram & cortical and subcortical segmentation
18
Including the anatomical information improves the overlap between automatic and manual tracts by 20%.
[189]
MLP & point-cloud-based networks
tractogram
42
@@
In addition to the streamline to be classified, 20 nearest neighbors and 500 streamlines randomly selected from the tractogram are used as input to the MLP. The latent representation computed by the MLP is used by the point cloud network to classify the streamline.
[190]
Bayesian model
tractogram and anatomical MRI
14
This model addresses the neonatal age range and is in part based on [173].
@@
  • A large class of automatic tract delineation methods is based on processing of whole-brain tractograms [199, 200, 201, 188, 202]. There are two common approaches used by these methods. (1) One set of methods compare individual streamlines with a predefined set of fibers in an atlas and assign the streamline to a specific tract [203, 199, 200]. (2) Other methods cluster the streamlines based on some measure of pair-wise similarity [204, 182]. There exist various techniques for implementing the clustering, for example using normalized cuts [205]. Both these sub-classes of methods can be computationally expensive, especially for larger number of streamlines and higher image resolution. Consequently, much effort has been directed towards developing fast streamline clustering methods [204, 206]. Some techniques additionally take into account the location of the streamlines relative to anatomical landmarks in the brain [188, 202, 207, 208]. Most clustering-based methods aim at finding streamlines that have similar geometric shapes and spatial extents. There are also methods that cluster streamlines based on their start/end points [209]. However, methods that use streamline start/end points may not perform as well as streamline clustering methods [210].
@@
  • Overall, methods that use anatomical information achieve better results than methods that rely on streamline shape alone. Such information can be encoded and incorporated in the tract segmentation computation in various ways. For example, Maddah et al. [207] used the anatomical information provided by an atlas of white matter tracts, while Tunc et al. [211, 212] used the connectivity signature of the streamlines. Yendiki [173] has proposed a method that uses the tractography results computed from the dMRI signal and the anatomical tissue segmentation information for automatic tract delineation. Theirs is a probabilistic method, where prior probabilities of the tracts is computed from a set of manually-segmented training subjects. This method, trained with data from healthy subjects, performed well on schizophrenia patients. This method has been extended to neonatal brain using a dedicated image processing pipeline to address the challenges of early brain data [190]. This extended method also showed good generalizability to data from term- and preterm-born neonates imaged with different scanners and acquisition protocols.
@@
    Sydnor et al. [213] critically compared a manual ROI-based method [214] and two automatic tractography-based tract segmentation techniques [209, 182] in terms of various measures of accuracy and reproducibility. They concluded that although clustering methods were overall better, they were not consistently superior and that an optimal method might have to rely on a combination of approaches. In a rather similar work, O’Donnell et al. [215] compared streamline clustering methods with techniques that clustered streamlines based on their start and end points in the cortex. They highlighted the pros and cons of each class of methods and advocated for hybrid methods that could benefit from the strengths of both classes of techniques. Similarly, Chekir at al. [216] and Xu et al. [217] point out the limitations of the purely clustering-based methods and methods that are based on anatomical priors. They propose methods that synergistically combine the advantages of the two strategies.
@@
  • A growing number or works have applied deep learning methods on whole-brain tractograms to extract/cluster individual tracts. The proposed methods have followed very different designs. A typical recent example is Deep White Matter Analysis (DeepWMA) [175]. DeepWMA first computes a set of feature descriptors to represent the spatial coordinate location of points along each streamline. To render the feature descriptor independent of the orientation/order of the tracing of the streamline, the coordinates are also flipped and added to the descriptor via concatenation. The resulting 1D representation is then repeated row-wise to create a 2D feature map that a CNN can efficiently process. A CNN is trained to classify the streamlines based on the feature maps. In order to enable their model to identify false positive streamlines that are very common in tractography results, they include an extra class label to represent those streamlines. DeepBundle [176], on the other hand, adopts a representation in terms of the coordinates of the points along streamlines as functions defined on a graph. Using the coordinates as input, a graph CNN is applied to extract geometric features and to classify the streamlines. Hence, DeepBundle sidesteps the initial manual feature engineering that is common in many tractography-based methods. The graph CNN, itself, includes a series of graph convolution and graph pooling layers. TRAFIC [177] computes streamline curvature, torsion, and Euclidean distances to selected anatomical landmarks for each point on the streamline. Fibernet [179, 180], on the other hand, represents the streamlines using an iso-surface map of the brain. At least one study has observed that using streamline coordinates as input leads to more accurate tract segmentation than features such as curvature and torsion [181]. TractCloud [189] uses the streamline to be classified, its nearest neighbors, and a set of randomly selected streamlines from the whole-brain tractogram as input to an MLP to compute a latent representation, which is then used by a point cloud network to classify the streamline.

### v2

@@ IV-F1 Method design
An increasingly popular class of models in tractography is recurrent neural networks (RNNs) [234, 233, 235, 236, 239]. Poulin et al. [234] trained MLPs and Gated Recurrent Units (GRUs) [240] to map the dMRI signal to the streamline directions. The input to their model was the b0-normalized signal, resampled to a fixed spherical grid with 100 directions. The information from the preceding tracing steps was passed to the current step via the hidden state vector. Experiments showed that the GRU-based method outperformed 96 competing techniques on a simulated dataset and achieved expert-level results on a scan from the HCP dataset. Benou et al. [235] used an RNN to estimate a probabilistic FOD, which could then be used for deterministic or probabilistic tractography. The streamline tracing was stopped when the model predicted a termination label or when the entropy of the predicted FOD was below a threshold. This method showed competitive performance compared with classical methods and machine learning techniques, especially in terms of false positive rate and ability to reconstruct different fiber bundles.
@@
The majority of the proposed methods have followed a supervised learning strategy. The main difficulty with these approaches is obtaining rich data with reliable ground truth, discussed further in the next sub-section. Given the difficulty of obtaining training data, there is a growing interest in reinforcement learning techniques [241, 237, 242]. Theberge et al. developed the first deep reinforcement learning-based tractography method [237]. In their method, the state is derived from the diffusion signal while the reward is computed from the FOD peaks and the direction of the streamline. The model input consists of SH-represented FOD at the current voxel and its six neighbors. Additionally, the past four streamline directions are included as extra inputs to the model. The reward function is based on local information. Specifically, it promotes the closeness of the predicted streamline tracing step with the major FOD peaks and encourages streamline smoothness. Experiments show that the reinforcement learning method is competitive with supervised learning methods while also having better generalizability to unseen datasets. However, a true reinforcement learning method requires the knowledge of streamline start and end points, which is unavailable for in-vivo data.
@@
IV-F2 Training data
One approach to obtaining ground truth tractography data is via elaborate manual editing of automatically generated tractograms by experts [244]. However, even for experts it is often impossible to resolve ambiguous situations and this practice suffers from high inter-observer variability [245, 246]. Physical phantoms represent an alternative [247, 248]. Poupon et al. have developed the FiberCup phantom to simulate a coronal slice of the brain at the level of corticospinal tracts [249, 250]. This phantom includes a rich set of crossing, fanning, and kissing tracts as well as U-fibers. The main problem with this approach is that no phantom can represent the complexity and inter-subject variability of the human brain. Lack of reliable data for model development and validation has remained one of the main stumbling blocks for advancing machine learning-based tractography, similar to other applications discussed in this paper. A notable recent database that has been made publicly available is TractoInferno [251]. It includes multi-scanner data acquired with different protocols. In addition to the processed dMRI data, TractoInferno includes 30 white matter bundles reconstructed using an ensemble of 4 different tractography approaches followed by automatic and manual quality control.
IV-F3 Evaluations and main results
Machine learning methods have been effective in reconstructing various tracts with low false positive rates, reducing the systematic errors and biases that are prevalent with classical tractography methods, and improving generalizability to unseen data [56, 18]. Several reasons have been put forward to explain the success of machine learning methods. First, they avoid making rigid, ad-hoc, and possibly sub-optimal modeling assumptions or streamline propagation and stopping rules that are common in standard methods [56, 18, 238]. Some of the proposed machine learning methods use the dMRI data (rather than diffusion tensor or FOD estimates) as input, thereby sidestepping the unavoidable errors in fiber orientation estimation [56, 235, 59]. Moreover, they have the potential to learn tissue probabilities from dMRI data, rather than relying on the information provided by a registered anatomical MRI or ambiguous parameters such as FA. Machine learning methods can also learn the noise and artifacts from data, instead of imposing simplified noise models and assuming artifact-free scans [56]. Furthermore, these methods can incorporate estimation uncertainty into streamline tractography in a systematic way [238]. Machine learning techniques are also by far more flexible in incorporating non-local information, such as anatomical context, which can greatly improve tractography results [18]. An important piece of information that can be easily incorporated into most machine learning models is the direction of preceding streamline tracing steps. The direction of the last step indicates the streamline slope and the direction of the last two steps indicate its curvature. This information has been useful in improving tractography accuracy [234]. Incorporation of the direction of preceding steps can also improve the tractography in voxels with crossing fibers and bottleneck regions [237, 234].
IV-F4 Tractography post-processing
Machine learning methods have also been used to enhance the generated tractograms via post-processing operations such as identification of false positives or anatomically implausible streamlines. Deep learning has shown great success at this task as well. An example of a supervised method is the work of Astolfi [252], where an existing method [253] is used to generate labels on training data based on anatomical priors. A limitation of supervised approaches is that it is hard to obtain training data that can adequately represent the complete range of invalid streamlines [254]. Unsupervised methods sidestep this limitation. An example of unsupervised techniques is the auto-encoder method proposed by Legarreta [255]. This method uses unlabeled whole-brain tractograms to learn a low-dimensional representation of valid streamlines, which can then be used to detect spurious or anatomically invalid streamlines in the test data. Both supervised and unsupervised approaches have reported remarkable success.
IV-G Delineation of white matter tracts
@@
The brain white matter is organized into distinct tracts, which consist of bundles of myelinated axons that connect different brain regions such as the cerebral cortex and the deep gray matter [256, 257, 258]. Although they are tightly packed and often cross one another, each tract has a different function and connects different regions of the brain. Accurate segmentation of these tracts is needed in clinical studies and medical research. For example, in surgical planning one needs to know the precise extent of the individual tracts in order to assess the risk of damage to specific neurocognitive functions that may result from surgical removal of brain tissue [259, 224]. Furthermore, changes in the tissue microstructure on specific tracts can be associated with brain development and disorders [260, 261, 262].
@@
Diffusion MRI is the only non-invasive method that can accurately delineate white matter tracts [258]. Individual tracts may be extracted from whole-brain tractograms by specifying inclusion and exclusion regions of interest (ROIs). This process, which is usually referred to as “virtual dissection”, is time-consuming, requires substantial expertise, and suffers from high inter-observer variability and low reproducibility [246, 263].
Several prior works have aimed at automating the virtual dissection process by learning to compute the inclusion/exclusion ROIs [264, 265, 266]. Moreover, there has been much effort to develop fully automatic methods for tract delineation. These methods can be divided into two categories: (1) methods that start by performing streamline tractography; (2) methods that do not depend on tractography. We have summarized some of the works in these two classes of methods, respectively, in Tables VII and VIII. Given the large number of published methods, these tables only list a selection of more recent methods with a focus on deep learning techniques. Below, we describe our main findings from our study of these works.
@@
TABLE VII: A summary of selected tractography-based methods for automatic delineation/segmentation of white matter tracts.
  Reference Model Input Number of tracts Methodology and results
@@
  Deep white matter analysis (DeepWMA) [267] CNN Tractogram 54 The method is based on hand-crafted features describing the streamline geometry. A CNN is applied to classify the streamlines based on the feature vectors. The method shows high accuracy across the human lifespan from neonatal to old age and it works well on brains with gross pathologies such as tumors.
DeepBundle [214] Graph CNN Tractogram 12 A graph CNN is used to classify streamlines based on their geometry. Only HCP data is used in this work. The method shows higher accuracy than standard methods, especially for small tracts such as fornix and commissure anterior.
BrainSegNet [268] Bidirectional LSTM Tractogram 8 The method works in two steps, in the first step classifying the white matter versus the gray matter and in the second step segmenting the tracts. The method achieves streamline classification accuracy of $>96\%$ and recall of $>73\%$, but it is tested on three subjects only.
FiberNET [269] CNN Tractogram 17 The method uses a harmonic function to compute level-sets based on a brain shape-center, thereby re-parameterizing the streamlines in a manner that is consistent across subjects. A CNN classifies the streamlines based on this representation. The method achieves low false positive rates.
Xu et al. [270] CNN Tractogram 64 The focus of this work was on detecting eloquent white matter tracts for epilepsy surgery. Different CNN architectures and loss functions were investigated. The models were tested on healthy children and children with epilepsy. A deep CNN with attention mechanism achieved the highest F1 score of 0.95.
@@
Spectral clustering [271] Spectral embedding clustering Tractogram 10 The method achieves high tract segmentation accuracy and can be useful for cross-subject white matter studies.
@@
Dayan et al. [272] Restricted Boltzmann machines (RBM) Tractogram 61 The method offers memory-efficient unsupervised detection of anatomically relevant tracts without the need to explicitly define the tracts. It achieves Dice Similarity Coefficients ranging from under 0.30 to above 0.90 on different tracts.
@@
TRACULA [265] Bayesian model Tractogram and anatomical MRI 18 The model trained on healthy subjects worked well on schizophrenia patients.
FS2Net [273] LSTM Tractogram 8 The authors claim that, because their method relies on streamline geometry, it is particularly useful in situations where the test brains are arbitrarily oriented. They report accuracies above 0.90 but only on three subjects.
@@
Deep Fiber Clustering [274] Siamese Graph convolutional neural network Tractogram and gray matter parcellation Unclear The method produces accurate delineation of tracts across age and gender for healthy as well as diseased brains.
Superficial White Matter Analysis (SupWMA) [275] MLP Tractogram 198 This method is based on a point cloud-based representation of the streamlines and contrastive supervised learning. Experimental results on six datasets across the human lifespan show that this approach can accurately extract superficial white matter tracts.
Ugurlu et al. [254] Ensemble of MLPs Tractogram and FOD map 9 This paper presents a way of synthesizing a rich set of invalid streamlines to enable effective training of the model to reject such streamlines in test data. The method is only validated on HCP data.
AnatomiCuts [276] Hierarchical clustering Tractogram & cortical and subcortical segmentation 18 This method advocates for leveraging the position of the streamline with respect to cortical and sub-cortical landmarks. Experiments on HCP data show that this approach improves the overlap between automatic and manual tracts by 20%.
@@
TractCloud [277] MLP & Point-cloud-based networks Tractogram 42 In addition to the streamline to be classified, 20 nearest neighbors and 500 streamlines randomly selected from the tractogram are used as input to the MLP. The latent representation computed by the MLP is used by the point cloud network to classify the streamline. The method is validated on five datasets.
TRACULInA [278] Bayesian model Tractogram and anatomical MRI 14 This model addresses the neonatal age range and is in part based on TRACULA [265]. It is successfully validated on out-of-distribution data such as prematurely-born neonates.
IV-G1 Methods that rely on tractography
A large class of automatic tract delineation methods is based on processing of whole-brain tractograms [279, 280, 281, 276, 282, 255]. There are two common approaches used by these methods.
  1. 1.
    One set of methods compare individual streamlines with a predefined set of fibers in an atlas and assign each streamline to a specific tract in the atlas [283, 279, 280].
  2. 2.
    Another set of methods cluster the streamlines based on some measure of pair-wise similarity [284, 271]. There exist various techniques for implementing the clustering, for example using normalized cuts [285] or k-nearest-neighbors [286].
@@ IV-G1 Methods that rely on tractography
Both these sub-classes of methods can be computationally expensive, especially for larger number of streamlines and higher image resolution. Consequently, much effort has been directed towards developing fast streamline clustering methods [287, 284, 288]. Some techniques additionally take into account the location of the streamlines relative to anatomical landmarks in the brain [276, 282, 289, 290]. Most clustering-based methods aim at finding streamlines that have similar geometric shapes and spatial extents. There are also methods that cluster streamlines based on their start/end points [291]. However, methods that use streamline start/end points may not perform as well as streamline clustering techniques [292].
@@
Overall, methods that use anatomical information achieve better results than methods that rely on streamline shape alone. Such information can be encoded and incorporated in the tract segmentation computation in various ways. For example, Maddah et al. [289] used the anatomical information provided by an atlas of white matter tracts, while Tunc et al. [293, 294] used the connectivity signature of the streamlines. Yendiki [265] has proposed a method that uses the tractography results computed from the dMRI signal and the anatomical tissue segmentation information for automatic tract delineation. Theirs is a probabilistic method, where prior probabilities of the tracts are computed from a set of manually-segmented training subjects. This method, trained with data from healthy subjects, performed well on schizophrenia patients. It has been extended to neonatal brain using a dedicated image processing pipeline to address the challenges of early brain data [278]. This extended method also showed good generalizability to data from term- and preterm-born neonates imaged with different scanners and acquisition protocols.
@@
Sydnor et al. [295] critically compared a manual ROI-based method [296] and two automatic tractography-based tract segmentation techniques [291, 271] in terms of various measures of accuracy and reproducibility. The results showed that although clustering methods were overall better, they were not consistently superior and that an optimal method might have to rely on a combination of the two approaches. In a rather similar work, O’Donnell et al. [297] compared streamline clustering methods with techniques that clustered streamlines based on their start and end points in the cortex. They highlighted the pros and cons of each class of methods and advocated for hybrid methods that could benefit from the strengths of both classes of techniques. Similarly, Chekir at al. [298] and Xu et al. [299] point out the limitations of the purely clustering-based methods and methods that are based on anatomical priors. They propose methods that synergistically combine the advantages of the two strategies.
@@
A growing number or works have applied deep learning methods on whole-brain tractograms to extract/cluster individual tracts. These methods have followed very different designs. Given the high flexibility in designing deep learning methods, some of them cannot be categorized under any of the two broad classes mentioned above. A typical recent example is Deep White Matter Analysis (DeepWMA) [267]. DeepWMA first computes a set of feature descriptors to represent the spatial coordinate location of points along each streamline. In order to render the feature descriptor independent of the orientation/order of streamline, the coordinates are also flipped and added to the descriptor via concatenation. The resulting 1D representation is then repeated row-wise to create a 2D feature map that a CNN can efficiently process. A CNN is trained to classify the streamlines based on the feature maps. In order to enable their model to identify false positive streamlines, they include an extra class label to represent those streamlines. In DeepBundle [214], on the other hand, streamlines are represented in terms of the point coordinates and a graph CNN is trained to extract geometric features to classify the streamlines. Hence, DeepBundle sidesteps the manual feature engineering that is used in some tractography-based methods. The graph CNN includes a series of graph convolution and graph pooling layers. TRAFIC [300] computes streamline curvature, torsion, and Euclidean distance to anatomical landmarks for each point on the streamline. Fibernet [301, 269], on the other hand, represents the streamlines using an iso-surface map of the brain. At least one study has observed that using streamline coordinates as input leads to more accurate tract segmentation than features such as curvature and torsion [270]. TractCloud [277] uses the streamline to be classified, its nearest neighbors, and a set of randomly selected streamlines from the whole-brain tractogram as input to an MLP to compute a latent representation, which is then used by a point cloud network to classify the streamline.
IV-G2 Methods that avoid tractography